# Building sets, Chow rings, and their Hilbert series

### v1

@@ 6.2. Koszulity, gamma-positivity, and real-rootedness
A result by Coron [Cor25b, Theorem 3.12] states that if the lattice of flats $\mathcal{L}(\mathsf{M})$ of a matroid is supersolvable, then the Chow ring $D(\mathsf{M},\mathcal{G}_{\min})$ is Koszul. This generalizes a result by Dotsenko [Dot22] for the braid matroid $\mathsf{K}_{n}$. On the other hand, Mastroeni and McCullough showed that for all matroids $\mathsf{M}$, the ring $D(\mathsf{M},\mathcal{G}_{\max})$ is Koszul.
As is explained in [FMSV24, Section 5.4], whenever the degree of the Hilbert series of a graded Artinian Koszul algebra is not more than $4$, then it has only real zeros. Furthermore, by [FMSV24, Theorem 1.8], for any matroid $\mathsf{M}$ we have that $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}_{\max}}(x)$ is $\gamma$-positive.
@@
We note that the real-rootedness question for $\mathrm{H}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)$ was posed by Aluffi, Chen, and Marcolli in [ACM24, Conjecture 1], while in this special case, $\gamma$-positivity follows directly from Proposition 4.6 (see [ACM24, Theorem 1.2]).

### v2

@@ 6.2. Koszulity, gamma-positivity, and real-rootedness
A result by Coron [13, Theorem 3.12] states that if the lattice of flats $\mathcal{L}(\mathsf{M})$ of a matroid is supersolvable, then the Chow ring $D(\mathsf{M},\mathcal{G}_{\min})$ is Koszul. This generalizes a result by Dotsenko [16] for the braid matroid $\mathsf{K}_{n}$. On the other hand, Mastroeni and McCullough showed that for all matroids $\mathsf{M}$, the ring $D(\mathsf{M},\mathcal{G}_{\max})$ is Koszul.
As is explained in [21, Section 5.4], whenever the degree of the Hilbert series of a graded Artinian Koszul algebra is not more than $4$, then it has only real zeros. Furthermore, by [21, Theorem 1.8], for any matroid $\mathsf{M}$ we have that $\mathrm{H}_{\mathsf{M}}^{\mathcal{G}_{\max}}(x)$ is $\gamma$-positive.
@@
We note that the real-rootedness question for $\mathrm{H}_{\mathsf{K}_{n}}^{\mathcal{G}_{\min}}(x)$ was posed by Aluffi, Chen, and Marcolli in [2, Conjecture 1], while in this special case, $\gamma$-positivity follows directly from Proposition 4.6 (see [2, Theorem 1.2]).